# Residual Prior-driven Frequency-aware Network for Image Fusion

### v1

@@ Abstract.
Image fusion aims to integrate complementary information across modalities to generate high-quality fused images, thereby enhancing the performance of high-level vision tasks. While global spatial modeling mechanisms show promising results, constructing long-range feature dependencies in the spatial domain incurs substantial computational costs. Additionally, the absence of ground-truth exacerbates the difficulty of capturing complementary features effectively. To tackle these challenges, we propose a Residual Prior-driven Frequency-aware Network, termed as RPFNet. Specifically, RPFNet employs a dual-branch feature extraction framework: the Residual Prior Module (RPM) extracts modality-specific difference information from residual maps, thereby providing complementary priors for fusion; the Frequency Domain Fusion Module (FDFM) achieves efficient global feature modeling and integration through frequency-domain convolution. Additionally, the Cross Promotion Module (CPM) enhances the synergistic perception of local details and global structures through bidirectional feature interaction. During training, we incorporate an auxiliary decoder and saliency structure loss to strengthen the model’s sensitivity to modality-specific differences. Furthermore, a combination of adaptive weight-based frequency contrastive loss and SSIM loss effectively constrains the solution space, facilitating the joint capture of local details and global features while ensuring the retention of complementary information. Extensive experiments validate the fusion performance of RPFNet, which effectively integrates discriminative features, enhances texture details and salient objects, and can effectively facilitate the deployment of the high-level vision task.
This paper has been accepted by ACM MM 2025, and the code has been released publicly.
Multi-modal Image Fusion, Residual Prior, Frequency Domain, Contrastive Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer vision
@@ 1. Introduction
Image fusion serves as an enhancement technique aimed at integrating complementary scene information captured by multiple sensors, generating information-rich fused images that overcome the limitations of single sensors and facilitate the deployment of high-level vision tasks (2, ; 3, ). Infrared and visible image fusion (IVIF) represents a critical research domain that combines the thermal information from infrared (IR) images with the rich textural details of visible (VIS) images, producing comprehensive scene representations that mitigate the illumination sensitivity of visible images and the noise susceptibility and low resolution of infrared images (1, ). This comprehensive information integration provides solutions for a wide range of scenarios and requirements (4, ; 5, ; 6, ; 7, ; 8, ; 9, ; 10, ; 11, ; 12, ).
@@
With the advancement of deep learning (DL) techniques, IVIF has gained more effective guidance, particularly when addressing challenges posed by various environments and conditions. A typical pipeline that utilizes the nonlinear capabilities of CNNs to learn fusion mappings has shown promising results due to its strong generalization ability. However, CNNs have limitations in their ability to model long-range dependencies, as they only capture interactions within their receptive fields, which presents challenges when balancing texture and saliency objectives. To enhance fusion performance, Transformer-based architectures were introduced, utilizing self-attention mechanisms to capture global contextual interactions and model long-range dependencies. Nevertheless, these methods incur high computational complexity and parameter counts, with self-attention computation costs growing quadratically when processing complex images, limiting their application in resource-constrained environments. Figure 1 (a) and (b) provide intuitive examples. Furthermore, the lack of direct ground truth complicates the learning of fusion mappings, as existing methods typically rely on proxy losses. In this context, effectively capturing and integrating complementary information from different modalities remains challenging.
@@
To tackle these challenges, we propose a Residual Prior-driven Frequency-aware Network (RPFNet) for image fusion based on the following key observations: 1) Residual maps between different modalities contain significant textural features that provide crucial prior guidance for fusion (50, ; 51, ); 2) Processing features in the frequency domain offers more efficient global information acquisition compared to modeling long-range dependencies in the spatial domain (See Figure 1 (c)); 3) Coupled convolution across the spatial and frequency domains can achieve collaborative perception of local-global features and enhance the feature representation ability of the model (5, ; 6, ; 30, ). Based on these observations, we design a dual-branch feature extraction architecture incorporating a Residual Prior Module (RPM) and a Frequency Domain Fusion Module (FDFM). The RPM captures fine-grained multi-modal complementary features by mining multi-scale information from residual maps, while the FDFM maps features to the frequency domain through Fourier transformation, achieving efficient global information integration at lower computational cost than spatial domain modeling of long-range dependencies. In particular, we develop a Cross-Promotion Module (CPM) for bidirectional feature enhancement: On one hand, CPM introduces the residual prior information extracted by RPM into FDFM, providing critical modal difference guidance for frequency domain fusion, allowing it to focus on complementary information between modalities. On the other hand, CPM simultaneously feeds back fused features generated by FDFM to RPM, promoting more refined difference feature perception in the residual prior module. During the training process, we introduce an auxiliary decoder coupled with a specifically designed saliency structural loss $\mathcal{L}_{SS}$ to further enhance the model’s capability to perceive modality-specific differences in residual mappings. Additionally, we propose an adaptive weight-based frequency contrastive fidelity term $\mathcal{L}_{c}$, which utilizes contrastive learning in the frequency domain to shrink the solution space while driving the model to achieve global content awareness. This, combined with the locally-operating SSIM function (39, ), equips the model with the ability to capture both local and global features, ensuring the effective preservation of complementary features. Extensive experimental validation has been conducted to demonstrate the proposed method’s effectiveness in fusion performance and high-level vision task applications. In summary, our main contributions are as follows:
@@ 2. Related work
Frequency domain learning. Fast Fourier Convolution (24, ) has demonstrated impressive performance by transferring features from the spatial domain to the frequency domain, enabling global-to-global feature mapping. This has inspired researchers to seamlessly integrate a series of frequency domain techniques into deep learning frameworks, facilitating the learning of non-local features (24, ; 25, ; 26, ; 29, ). Kong et al. (27, ) proposed a frequency-domain attention solver to reduce model complexity. Chen et al. (28, ) investigated dilated convolution from a frequency perspective and introduced frequency-adaptive convolution to enhance the capabilities of dilated operations. Furthermore, the inherent resilience of frequency domain features to local variations and noise has led to remarkable success in various image restoration tasks, including denoising (32, ; 33, ), deraining (30, ; 31, ), and deblurring (27, ). Following these advances, Zheng et al. (5, ) explored IVIF in the frequency domain, effectively mining complementary features from source images through the synergistic operation of spatial and frequency domains. Zheng et al. (6, ) also leveraged higher-order interactions within the frequency domain to capture the synergy between spatial fine-grained details and global statistics across multiple modalities. Building on this insights, Zhou et al. (34, ) further investigated the synergistic effects between spatial fine-grained features and global statistics across modalities, proposing a novel paradigm for collaborative high-order interaction.
Contrastive learning for image fusion. Contrastive learning has emerged as a classical paradigm in image fusion for its ability to capture implicit feature relationships without ground truth by constructing positive and negative sample pairs to drive the learning of generalizable features (4, ; 35, ; 36, ; 37, ; 52, ). Wang et al. (4, ) constructed positive and negative samples based on the similarity and divergence between source features, enabling better extraction of complementary modal features. Liu et al. (36, ) developed a coupled contrastive learning framework utilizing manual annotations to balance the representation of salient objects and texture details. Zhu et al. (37, ) proposed a patch-based approach for constructing positive-negative pairs to preserve source image features. Wang et al. (35, ) introduced unpaired high-quality images as positive samples to achieve degradation-resistant feature fusion.
@@ 4.1. Experimental Setting
To evaluate the effectiveness and generalizability of our method, we compared it with 9 SOTA methods: LRRNet (9, ), CS2Fusion (4, ), Text-IF (7, ), FISCNet (5, ), TarDAL (13, ), IGNet (8, ), SHIP (6, ), Conti-Fuse (10, ), and A2RNet (14, ). We employed 6 metrics: three non-reference metrics (EN, SF, SD) and three reference metrics (CC, SCD, and VIFF). For all these metrics, higher values indicate better fusion performance (2, ; 3, ).
@@ 4.2.3. Fusion result on the test dataset
Figure 5 and Table 4 present the qualitative and quantitative results of our proposed method on the test dataset. Our method effectively preserves complementary features from source images through residual prior guidance and global modeling in the frequency domain. Additionally, the adaptive weight-based frequency domain contrastive learning enables our model to better differentiate textural details and high-contrast features. This characteristic allows our method to penetrate through smoke and overexposed environments to better represent scene information—a comprehensive capability that other SOTA methods struggle to achieve. Notably, while frequency domain modeling approaches like SHIP and FISCNet are effective, they focus primarily on modeling complementary features while neglecting the guidance of residual priors, resulting in a gap in discriminative feature expression compared to our method. Quantitatively, our method’s dominant position further confirms its effectiveness. High scores on EN, SD, and SF metrics demonstrate our method’s significant capability in integrating textural details, while performance on CC, SCD, and VIFF metrics indicates superior integration of complementary information from source images, highlighting its ability to transfer rich information from source images to the fused result. Although our method shows a slight deficiency in SCD scores on MSRS, from an overall perspective, it still demonstrates robustness and generalizability.
@@
4.3. Extended Experiments
4.3.1. Evaluation with Power Spectral Density
4.3.2. Extension to object detection
@@ 5. Conclusion
This research proposes a Residual Prior-driven Frequency-domain Aware Fusion Network that efficiently integrates global information through frequency-domain convolution while leveraging saliency structural priors from residual mapping to preserve complementary cross-modal information without compromising computational efficiency. The meticulously designed CPM employs a bidirectional closed-loop architecture that effectively retains both local details and salient features. During training, an adaptive-weight-based frequency-domain contrastive loss facilitates the preservation of complementary features from a global perspective, while the saliency structure loss utilizes the auxiliary decoder to enable the model to capture modality-specific differences. These effective design elements enable the proposed method to simultaneously highlight salient objects while preserving fine-grained details, consequently enhancing performance across high-level vision tasks, as demonstrated through extensive experimental validation.

### v2

@@ Abstract.
Image fusion aims to integrate complementary information across modalities to generate high-quality fused images, thereby enhancing the performance of high-level vision tasks. While global spatial modeling mechanisms show promising results, constructing long-range feature dependencies in the spatial domain incurs substantial computational costs. Additionally, the absence of ground-truth exacerbates the difficulty of capturing complementary features effectively. To tackle these challenges, we propose a Residual Prior-driven Frequency-aware Network, termed as RPFNet. Specifically, RPFNet employs a dual-branch feature extraction framework: the Residual Prior Module (RPM) extracts modality-specific difference information from residual maps, thereby providing complementary priors for fusion; the Frequency Domain Fusion Module (FDFM) achieves efficient global feature modeling and integration through frequency-domain convolution. Additionally, the Cross Promotion Module (CPM) enhances the synergistic perception of local details and global structures through bidirectional feature interaction. During training, we incorporate an auxiliary decoder and saliency structure loss to strengthen the model’s sensitivity to modality-specific differences. Furthermore, a combination of adaptive weight-based frequency contrastive loss and SSIM loss effectively constrains the solution space, facilitating the joint capture of local details and global features while ensuring the retention of complementary information. Extensive experiments validate the fusion performance of RPFNet, which effectively integrates discriminative features, enhances texture details and salient objects, and can effectively facilitate the deployment of the high-level vision task. The source code can be available at https://github.com/wang-x-1997/RPFNet.
This paper has been accepted by ACM MM 2025
Multi-modal Image Fusion; Residual Prior; Frequency Domain; Contrastive Learning
copyright: none
@@ 1. Introduction
Image fusion serves as an enhancement technique aimed at integrating complementary scene information captured by multiple sensors, generating information-rich fused images that overcome the limitations of single sensors and facilitate the deployment of high-level vision tasks (2, ; 3, ; 53, ; 35, ; 56, ). Infrared and visible image fusion (IVIF) represents a critical research domain that combines the thermal information from infrared (IR) images with the rich textural details of visible (VIS) images, producing comprehensive scene representations that mitigate the illumination sensitivity of visible images and the noise susceptibility and low resolution of infrared images. This comprehensive information integration provides solutions for a wide range of scenarios and requirements (4, ; 5, ; 6, ; 7, ; 8, ; 9, ; 10, ; 11, ; 12, ).
@@
With the advancement of deep learning (DL) techniques, IVIF has gained more effective guidance, particularly when addressing challenges posed by various environments and conditions. A typical pipeline that utilizes the nonlinear capabilities of CNNs to learn fusion maps has shown promising results due to its strong generalization ability. However, CNNs have limitations in their ability to model long-range dependencies, as they only capture interactions within their receptive fields, which presents challenges when balancing texture and saliency objectives. To enhance fusion performance, Transformer-based architectures were introduced, utilizing self-attention mechanisms to capture global contextual interactions and model long-range dependencies. Nevertheless, these methods incur high computational complexity and parameter counts, with self-attention computation costs growing quadratically when processing complex images, limiting their application in resource-constrained environments. Figure 1 (a) and (b) provide intuitive examples. Furthermore, the lack of direct ground truth complicates the learning of fusion maps, as existing methods typically rely on proxy losses. In this context, effectively capturing and integrating complementary information from different modalities remains challenging.
@@
To tackle these challenges, we propose a Residual Prior-driven Frequency-aware Network (RPFNet) based on the following key observations: 1) Residual maps between modalities contain critical textural priors (50, ; 51, ); 2) Frequency domain processing efficiently captures global information compared to spatial domain long-range modeling (Figure 1(c)); 3) Coupled convolution across spatial-frequency domains facilitates collaborative local-global feature perception (5, ; 6, ; 30, ). Specifically, RPFNet features a dual-branch architecture: a Residual Prior Module (RPM) capturing fine-grained multi-scale features from residual maps, and a Frequency Domain Fusion Module (FDFM) efficiently integrating global information via Fourier transformation. Moreover, a Cross-Promotion Module (CPM) enhances bidirectional interaction between RPM and FDFM. RPM provides residual priors to guide frequency fusion, while FDFM feedback refines residual feature perception. During training, an auxiliary decoder with a designed saliency structural loss $\mathcal{L}_{SS}$ strengthens modality-specific residual difference perception. An adaptive weight-based frequency contrastive fidelity term $\mathcal{L}_{c}$ shrinks solution space through frequency-domain contrastive learning, promoting global awareness. Coupled with local SSIM constraints (39, ), our model effectively captures complementary local-global features. Extensive experiments validate the method’s fusion performance and utility in high-level vision tasks. In summary, our contributions include:
@@ 2. Related work
Frequency Domain Learning. Recent advancements, such as Fast Fourier Convolution (24, ), leverage frequency domain techniques for efficient global feature map, inspiring various frequency-based methods (24, ; 29, ; 55, ; 57, ; 58, ). For instance, Kong et al. (27, ) designed frequency-domain attention to simplify models; Chen et al. (28, ) introduced frequency-adaptive convolution for enhanced dilated operations. Furthermore, frequency domain robustness has advanced image restoration tasks like denoising (32, ; 33, ), deraining (30, ; 31, ), and deblurring (27, ). In IVIF, Zheng et al. (5, ) mined complementary features using spatial-frequency synergies, and (6, ; 34, ) explored higher-order interactions to integrate fine-grained spatial details with global statistics.
Contrastive Learning for Image Fusion. Contrastive learning effectively captures implicit feature relationships without ground truth by constructing positive-negative sample pairs (4, ; 35, ; 36, ; 37, ; 52, ). Wang et al. (4, ) leveraged feature similarity-divergence to extract complementary modal features, while Liu et al. (36, ) employed annotations balancing salient and textural features. Zhu et al. (37, ) proposed patch-based positive-negative sampling for feature preservation, and Wang et al. (35, ) utilized unpaired high-quality images for degradation-resistant fusion.
@@ 4.1. Experimental Setting
To evaluate the effectiveness and generalizability of our method, we compared it with 9 SOTA methods: LRRNet (9, ), CS2Fusion (4, ), Text-IF (7, ), FISCNet (5, ), TarDAL (13, ), IGNet (8, ), SHIP (6, ), Conti-Fuse (10, ), and A2RNet (14, ). We employed 6 metrics: three non-reference metrics (EN, SF, SD) and three reference metrics (CC, SCD, and VIF). For all these metrics, higher values indicate better fusion performance (2, ; 3, ).
@@ 4.2.3. Fusion result on the test dataset
Figure 5 and Table 4 present the qualitative and quantitative results of our proposed method on the test dataset. Our method effectively preserves complementary features from source images through residual prior guidance and global modeling in the frequency domain. Additionally, the adaptive weight-based frequency domain contrastive learning enables our model to better differentiate textural details and high-contrast features. This characteristic allows our method to penetrate through smoke and overexposed environments to better represent scene information—a comprehensive capability that other SOTA methods struggle to achieve. Notably, while frequency domain modeling approaches like SHIP and FISCNet are effective, they focus primarily on modeling complementary features while neglecting the guidance of residual priors, resulting in a gap in discriminative feature expression compared to our method. Quantitatively, our method’s dominant position further confirms its effectiveness. High scores on EN, SD, and SF metrics demonstrate our method’s significant capability in integrating textural details, while performance on CC, SCD, and VIF metrics indicates superior integration of complementary information from source images, highlighting its ability to transfer rich information from source images to the fused result. Although our method shows a slight deficiency in SCD scores on MSRS, from an overall perspective, it still demonstrates robustness and generalizability.
4.3. Extended Experiments
4.3.1. Evaluation with Power Spectral Density
4.3.2. Extension to object detection
@@ 5. Conclusion
This research proposes a Residual Prior-driven Frequency-aware Network that integrates global information via frequency-domain convolution and leverages saliency structural priors from residual maps to preserve complementary cross-modal information. The CPM employs a bidirectional closed-loop architecture that retains both local details and salient features. During training, adaptive-weight-based frequency-domain contrastive loss preserves complementary features globally, while saliency structure loss uses an auxiliary decoder to capture modality-specific differences. These design elements enable the method to highlight salient objects while preserving fine-grained details, with extensive experiments validating its effectiveness on both static statistics and high-level vision tasks.
Acknowledgements.
This work was supported by National Natural Science Foundation of China under Grants No.62162065 and 61761045; Yunnan Fundamental Research Projects No.202201AT070167; Joint Special Project Research Foundation of Yunnan Province No.202401BF070001-023; Yunnan Province Visual and Cultural Innovation Team No.202505-AS350009.